OpenStudy (wasiqss):

Solve the DE by finding IF {(x^3)(y^2)-y}dx+{(x^2)(y^4)-x}dy=0

OpenStudy (wasiqss):

@JamesJ @satellite73 @amistre64 try it plz

OpenStudy (jamesj):

you had a question like this yesterday no?

OpenStudy (wasiqss):

yup but still im having difficulty figuring it out :|

OpenStudy (jamesj):

post a link to that other question will you.

OpenStudy (jamesj):

Where are you stuck?

OpenStudy (wasiqss):

w8 imma tell u in min.

OpenStudy (wasiqss):

james stuck in integration of\[\int\limits \left\{ (1)/\left[ (x^4)y-(x^2)(y^4) \right] \right\}\]

OpenStudy (wasiqss):

i can suppose x^2 to b X

OpenStudy (jamesj):

that is not an integral required in the solution. Where is that in the solution worked out on your other thread?

OpenStudy (wasiqss):

im doing this one and after applying rule 3, i have got this one , rest portion i have integrated and solved

OpenStudy (amistre64):

is this an exact diffy q by chance?

OpenStudy (wasiqss):

yup amistre it is

OpenStudy (amistre64):

whew!! is was hoping it would be simpler than it looked lol

OpenStudy (amistre64):

\[(x^3y^2-y)\ dx+(x^2y^4-x)\ dy=0\] integrate one side and derive it down to the other; then compare notes

OpenStudy (wasiqss):

and the other integral that giving me headache is \[\int\limits \left\{ 1/((x^3)(y^2)-x(y^5)) \right\}\]

OpenStudy (wasiqss):

amistre we have to do only by finding integrating factor of it :/

OpenStudy (amistre64):

\[\int (x^3y^2-y)\ dx\ = \frac{1}{4}x^4y^2-yx+g(y)\] \[Dy[\frac{1}{4}x^4y^2-yx+g(y)]=\frac{1}{2}x^4y-x+g'(y)\]

OpenStudy (amistre64):

integrating factors eh ....

OpenStudy (jamesj):

wasiqss, read your other thread, you were giving the integrating factor, right there!!

OpenStudy (wasiqss):

james i cant get u

OpenStudy (amistre64):

id read the other thread them cause I got no clue how to do this with integrating factors :)

OpenStudy (wasiqss):

amistre so not right :(

OpenStudy (amistre64):

just playing around with it ... \[\frac{1}{2}x^4y-x+g'(y)=x^2y^4-x\] \[g'(y)=x^2y^4-x-\frac{1}{2}x^4y+x\] \[\int g'(y)dy=\int x^2y^4-\frac{1}{2}x^4y\ dy \] \[g(y) =\frac{1}{5}x^2y^5-\frac{1}{4}x^4y^2+C \] im sure i did something wrong in there :/

OpenStudy (wasiqss):

amistre u need some brain storming as to how to solve these de's :P

OpenStudy (amistre64):

i do :) i do :)

OpenStudy (wasiqss):

@TuringTest u gonna do?

OpenStudy (turingtest):

I've never solved a exact DE problem like this the way ash did on your other post If I were to try this I would just be trying to imitate his process, which you should be equally capable of ask me a specific question about some part of the problem and perhaps I could help you

OpenStudy (wasiqss):

ok help me in the integral

OpenStudy (turingtest):

which one?

OpenStudy (turingtest):

\[\int\limits \left\{ 1/((x^3)(y^2)-x(y^5)) \right\}\]this one?

OpenStudy (wasiqss):

yup

OpenStudy (turingtest):

\[\int{dx\over x^3y^2-xy^5}\]with respect to x, right?

OpenStudy (turingtest):

hello?

OpenStudy (turingtest):

@wasiqss

OpenStudy (wasiqss):

yup

OpenStudy (turingtest):

ok, factor out the bottom as much as possible, treating y as constant

OpenStudy (wasiqss):

w8 w8

OpenStudy (wasiqss):

ok now what

OpenStudy (wasiqss):

1/(xy^2)(x^2-y^3)

OpenStudy (turingtest):

partial fractions

OpenStudy (wasiqss):

cause of the y's i get confused in partial fraction , can u help me a bit in that

OpenStudy (turingtest):

you can pull 1/y^2 out of the integral first too

OpenStudy (wasiqss):

so what remains is 1/(x)(x^2-y^3)

OpenStudy (turingtest):

yes, and you can do either A/x+(Bx+C)/(x^2-y^2) or I suppose use difference of squares and make it A/x+B/(x+y^(3/2))+C/(x-y^(3/2))

OpenStudy (wasiqss):

and the whole expression is equal to 1/y rite?

OpenStudy (turingtest):

no, once you multiple through by the LCM of the fractions, what's left over will be equal to 1

OpenStudy (wasiqss):

so this expression A/x+(Bx+C)/(x^2-y^2) is equal to 1?

OpenStudy (wasiqss):

and 1/y is constant that we will multiply after integration ?

OpenStudy (turingtest):

yes, we have to put the 1/y back in the end we are lust looking at the PF right now\[{A\over x}+{B\over x+y^{3/2}}+{C\over x-y^{3/2}}\]multiply through by \(x(x^2-y^3)\) to get rid of the fractions and what you get will be equal to 1

OpenStudy (turingtest):

\[A(x^2-y^3)+Bx(x-y^{3/2})+Cx(x+y^{3/2})=1\]

OpenStudy (turingtest):

I do not feel like solving for A, B, and C, so that is going to be your job ;)

OpenStudy (turingtest):

ok I'll give you A

OpenStudy (turingtest):

let x=0

OpenStudy (wasiqss):

thanks mate , know can u help with last integral

OpenStudy (wasiqss):

i did this one :)

OpenStudy (turingtest):

cool I have to go brother, but I'll be seeing you around I'm sure!

OpenStudy (wasiqss):

cant u help in another one

OpenStudy (turingtest):

Sorry, I have class to take and give today, so time for me to prepare next time friend

OpenStudy (wasiqss):

okay love u bro :)

OpenStudy (turingtest):

lol, 'love' is a strong word, but the platonic feeling is mutual XD

OpenStudy (wasiqss):

hehehe dont think too much about it cause i could not control my emotions because of u helping me :)

OpenStudy (wasiqss):

ash do the way u did other day

OpenStudy (wasiqss):

ash taking long time LOL

OpenStudy (ash2326):

Wasiqss do you want me to solve the complete or just the integration?

OpenStudy (wasiqss):

solving complete would be world of help to me!

OpenStudy (wasiqss):

u doing?

OpenStudy (ash2326):

Wait for some time. I'll have dinner and then I'd solve it here:)

OpenStudy (wasiqss):

ok thanks but remember to solve :)

OpenStudy (ash2326):

yeah

OpenStudy (wasiqss):

thanks :)

OpenStudy (ash2326):

I'll just solve the integral \[\int \frac{dx}{x^3y^2-xy^5} dx\] Let's multiply numerator and denominator by \(\large{x^{-3}}\) we'll get \[\int \frac{x^{-3}}{y^2-x^{-2}y^5}dx\] Let \[y^2-x^{-2}y^5=t\] so \[+2x^{-3}y^5dx=dt\] or \[x^{-3}dx=\frac{dt}{2y^5}dt\] we'll get \[\int \frac{\frac{dt}{2y^5}}{t}dt\] or \[\int \frac{dt}{{2y^5}t}dt\] We get \[\frac{\ln t}{2y^5}\] substitute value of t y^2-x^{-2}y^5 Finally we get \[\frac{\ln ({y^2-x^{-2}y^5)}}{2y^5}\]

OpenStudy (wasiqss):

last integral man!! (1)/((x^4)y-(x^2)(y^4)

OpenStudy (ash2326):

We have \[\int \frac{1}{x^4 y-x^2 y^4}dx\] Now we can write this as \[\int \frac{1}{x^2y(x^2 - y^3)}dx\] Let \[x=y^{3/2} sec \theta \] \[dx=y^{3/2} \sec \theta \tan \theta d\theta\] We get \[\int \frac{y^{3/2} \sec \theta \tan \theta}{y^3 sec^2 \theta \times y(y^3 \sec^2 \theta - y^3)}d \theta\] W\frac{ e get \[y^{-11/2} \int \frac{\sec \theta \tan \theta} {\sec ^2 \theta \times \tan^2 \theta} d\theta \] we get \[y^{-11/2} \int \frac{cos^2 \theta}{\sin \theta} d\theta\] We get \[y^{-11/2} (\int (cosec \theta- \sin \theta )d \theta\] I think you can do now:)

OpenStudy (wasiqss):

ash any alternative way , i mean by completing squate of PF?

OpenStudy (ash2326):

We have \[\frac{1}{y} \int \frac{1}{x^2(x^2-y^3)}dx\] Now rewriting this as \[\frac{1}{y^4} \int \frac{y^3}{x^2(x^2-y^3)}dx\] \[\frac{1}{y^4} \int \frac{x^2-(x^2-y^3)}{x^2(x^2-y^3)}dx\] We get now \[\frac{1}{y^4} \int (\frac{1}{x^2-y^3}-\frac{1}{ x^2} )dx\] Now we get \[\frac{1}{y^4} \int (\frac{1}{x^2-(y^{3/2})^2}-\frac{1}{ x^2} )dx\] Can you solve it now @wasiqss?

OpenStudy (wasiqss):

man how u became so good at integration?

OpenStudy (turingtest):

@wasiqss that last integral is almost identical to the one I solved with you above (though I like ash's way better)